### v1

# Multimodal foundation world models for generalist embodied agents

Abstract
Learning generalist embodied agents, able to solve multitudes of tasks in different domains is a long-standing problem. Reinforcement learning (RL) is hard to scale up as it requires a complex reward design for each task. In contrast, language can specify tasks in a more natural way. Current foundation vision-language models (VLMs) generally require fine-tuning or other adaptations to be functional, due to the significant domain gap. However, the lack of multimodal data in such domains represents an obstacle toward developing foundation models for embodied applications. In this work, we overcome these problems by presenting multimodal foundation world models, able to connect and align the representation of foundation VLMs with the latent space of generative world models for RL, without any language annotations. The resulting agent learning framework, GenRL, allows one to specify tasks through vision and/or language prompts, ground them in the embodied domain’s dynamics, and learns the corresponding behaviors in imagination. As assessed through large-scale multi-task benchmarking, GenRL exhibits strong multi-task generalization performance in several locomotion and manipulation domains. Furthermore, by introducing a data-free RL strategy, it lays the groundwork for foundation model-based RL for generalist embodied agents.
Figure 1: Multimodal foundation world models connect and align the video-language space of a foundation model with the latent space of a generative world model for reinforcement learning, requiring vision-only data. Our GenRL framework turns visual and/or language prompts into latent targets and learns to realize the corresponding behaviors by training in the world model’s imagination.
Project website:mazpie.github.io/genrl
1 Introduction
Foundation models are large pre-trained models endowed with extensive knowledge of the world, which can be readily adapted for a given task [41]. These models have demonstrated extraordinary generalization capabilities in a wide range of vision [27, 42, 62] and language tasks [40, 18, 50, 11]. As we aim to extend this paradigm to embodied applications, where agents physically interact with objects and other agents in their environment, we require generalist agents that are capable of reasoning about these interactions and executing action sequences within these settings [57].
Reinforcement learning (RL) allows agents to learn complex behaviors from visual and/or proprioceptive inputs [17, 25, 26] by maximizing a specified reward function. Scaling up RL to multiple tasks and embodied environments remains challenging as designing reward functions is a complicated process, requiring expert knowledge and prone to errors which can lead to undesired behaviors [1]. Recent work has proposed the adoption of visual-language models (VLMs) to specify rewards for visual environments using language [4, 45], e.g. using the similarity score computed by CLIP [42] between an agent’s input images and text prompts. However, these approaches require fine-tuning of the VLM [37] or adaptation of the visual domain [45], to work reliably.
In most RL settings, we lack multimodal data to train or fine-tune domain-specific foundation models, due to the costs of labelling agents’ interactions and/or due to the intrinsic unsuitability of some embodied contexts to be converted into language. For instance, in robotics, it’s non-trivial to convert a language description of a task to the agent’s actions which are hardware-level controls, such as motor currents or joint torques. These difficulties make it hard to scale current techniques to large-scale generalization settings, leaving open the question:
How does one effectively leverage foundation models for generalization in embodied domains?
In this work, we present GenRL, a novel approach for training generalist agents from visual or language prompts, requiring no language annotations. GenRL learns multimodal foundation world models (MFWMs), where the joint embedding space of a foundation video-language model [54] is connected and aligned with the representation of a generative world model for RL [22], using unimodal vision-only data. The MFWM allows the specification of tasks by grounding language or visual prompts into the RL domain. Then, we introduce an RL objective for learning to accomplish the specified tasks in imagination [23], by matching the prompts in latent space.
Compared to previous work in world models and VLMs for RL, one emergent property of GenRL is the possibility to generalize to new tasks in a completely data-free manner. After training the MFWM, it possesses both strong priors over the dynamics of the environment, and large-scale multimodal knowledge. This combination enables the agent to interpret a large variety of task specifications and learn the corresponding behaviors. Thus, analogously to foundation models for vision and language, GenRL allows generalization to new tasks without additional data and lays the groundwork for foundation models in embodied RL domains [41].  [^1] The code, datasets and trained models will be made publicly available.
2 Preliminaries and background
Additional related works can be found in Appendix A.
Problem setting. The agent receives from the environment observations $x\in\mathcal{X}$ and interacts with it through actions $a\in\mathcal{A}$. The objective of the agent is to accomplish a certain task $\tau$, which can be specified either in the observation space $x_{\tau}$, e.g. through images or videos, or in language space $y_{\tau}$, where $\mathcal{Y}$ represents the space of all possible sentences. Crucially, compared to a standard RL setting, we do not assume that a reward signal is available to solve the task. When a reward function exists, it is instead used to evaluate the agent’s performance.
Generative world models for RL. In model-based RL, the optimization of the agent’s actions is done efficiently, by rolling out and scoring imaginary trajectories using a (learned) model of the environment’s dynamics. In recent years, this paradigm has grown successful thanks to the adoption of generative world models, which learn latent dynamics by self-predicting the agent’s inputs [22]. World models have shown impressive performance in vision-based environments [23], improving our ability to solve complex and open-ended tasks [25]. Generative world models have been successfully extended to many applications, such as exploration [48], skill learning [39], solving long-term memory tasks [47], and robotics [55].
Foundations models for RL. Large language models (LLMs) have been used for specifying behaviors using language [38, 28, 53], but this generally assumes the availability of a textual interface with the environment or that observations and/or actions can be translated to the language domain. The adoption of vision-language models (VLMs) reduces these assumptions, as it allows the evaluation of behaviors in the visual space. However. this approach has yet to show robust performance, as it generally requires fine-tuning of the VLM [4, 15], prompt hacking techniques [9] or visual modifications to the environment [4].
Vision-language generative modelling. Given the large success of image-language generative models [46], recent efforts in the community have focused on replicating and extending such success to the video domain, where the temporal dimension introduces new challenges, such as temporal consistency and increased computational costs [29, 3]. Video generative models are similar to world models for RL, with the difference that generation models outputs are typically not conditioned on actions, but rather conditioned on language[29] or on nothing at all (i.e. an unconditional model).
3 GenRL
(a) Connecting and aligning (Section 3.2)
(b) Learning task behavior (Section 3.3)
Figure 2: Overview of GenRL. The agent learns a multimodal foundation world model that connects and aligns (a) the representation of a foundation VLM with the latent states of a generative world model. Given a certain task prompt, (b) the model allows embedding the task and translating into targets in the latent dynamics space, which the agent can learn to achieve by using RL in imagination.
3.1 World models for RL
GenRL learns a task-agnostic world model representation by modelling the sequential dynamics of the environment in a compact discrete latent space $S$ [23, 25]. Latent states $s\in S$ are sampled from independent categorical distributions. The gradients for training the model are propagated through the sampling process with straight-through estimation [5].
The world model is made of the following components:
Encoder: $q_{\phi}(s_{t}|x_{t})$, Sequence model: $h_{t}=f_{\phi}(s_{t-1},a_{t-1},h_{t-1})$,
Decoder: $p_{\phi}(x_{t}|s_{t})$, Dynamics predictor: $p_{\phi}(s_{t}|h_{t})$,
trained with the loss:
$\mathcal{L}_{\phi}=\sum_{t}\underbrace{D_{\mathrm{KL}}\bigl{[}q_{\phi}(s_{t}|x_{t})\|p_{\phi}(s_{t}|s_{t-1},a_{t-1})]}_{\textrm{dyn loss}}-\underbrace{\mathbb{E}_{q_{\phi}(s_{t}|x_{t})}[\log p_{\phi}(x_{t}|s_{t})\bigr{]}}_{\textrm{recon loss}},$ (1)
where $p_{\phi}(s_{t}|s_{t-1},a_{t-1})$ is a shorthand for $p_{\phi}(s_{t}|f_{\phi}(s_{t-1},a_{t-1},h_{t-1}))$. The sequence model is implemented as a linear GRU cell [8]. Differently from recurrent state space models (RSSM; [24]), for our framework, encoder and decoder models are not conditioned on the information present in the sequence model. This ensures that the latent states only contain information about a single observation. We can then leverage the encoder as a probabilistic visual tokenizer, that is grounded in the target embodied environment.
3.2 Multimodal foundation world models
Multimodal VLMs are large pre-trained models that have the following components:
Vision $\begin{split}\textrm{Vision embedder:}\quad e^{(v)}=f^{(v)}_{\textrm{PT}}(x_{t:t+k}),\end{split}\quad\quad\begin{split}\textrm{Language embedder:}\quad e^{(l)}=f^{(l)}_{\textrm{PT}}(y),\end{split}$
where $x_{t:t+k}$ is a sequence of visual observations and $y$ is a text prompt. For video-language models, $k$ is generally a constant number of frames (e.g. $k\in\{4,8,16\}$ frames). Image-language models are a special case where $k=1$ as the vision embedder takes a single frame as an input. For our implementation, we adopt the InternVideo2 video-language model [54].
To connect the representation of the multimodal foundation VLM with the world model latent space, we instantiate two modules: a latent connector and a representation aligner:
$\begin{split}\textrm{Connector:}\quad p_{\psi}(s_{t:t+k}|e),\\ \textrm{Aligner:}\quad e^{(v)}=f_{\psi}(e^{(l)}),\end{split}\quad\quad\begin{split}\mathcal{L_{\textrm{conn}}}=\sum_{t}D_{\mathrm{KL}}\bigl{[}p_{\psi}(s_{t}|s_{t-1},e)\|\textrm{sg}(q_{\phi}(s_{t}|x_{t}))\bigr{]},\\ \mathcal{L_{\textrm{align}}}=\|e^{(v)}-f_{\psi}(e^{(l)})\|^{2}_{2},\end{split}$
where $\textrm{sg}(\cdot)$ indicates to stop gradients propagating.
The connector learns to predict the latent states of the world model from embeddings in the VLM’s representation space. The connector’s objective consist of minimizing the KL divergence between its predictions and the world model’s encoder distribution. While more expressive architectures, such as transformers [52] or state-space models [20] could be adopted, we opt for a simpler GRU-based architecture for video modelling. This way, we keep the method simple and the architecture of the connector is symmetric with respect to the world model’s components.
Multimodal VLMs trained with contrastive learning exhibit a multimodality gap [33], where the spherical embeddings of different modalities are not aligned. The role of the aligner is to reduce this multimodality gap, by projecting text embeddings into their corresponding visual embeddings. Given a dataset of vision-language data, this projective function can be learned.
In embodied domains vision-language data is typically unavailable. Instead, we leverage the idea that language embeddings can be treated as a ‘corrupted’ version of their vision counterparts [61, 63]. This allows us to approximate the language embedding with the corresponding vision embedding: $e^{(l)}\approx e^{(v)}+\epsilon$.
Previous methods inject the noise into the embeddings when training the new module [61, 63], in our case the connector. However, we opt for training a separate aligner network. This allows us to train a noise-free connector, which has two main advantages: (i) it yields higher prediction accuracy for visual embedding inputs while maintaining a similar alignment for language embedding inputs; and (ii) it is more flexible; it’s easier to re-train/adapt for different noise levels, as it only requires re-training the aligner module, and its use can be avoided if unnecessary.
3.3 Learning specified task behaviors in imagination
World models can be used to imagine trajectories in latent space, using the dynamics model. This allows us to train behavior policies in a model-based RL fashion [23]. Given a task specified through a visual or language prompt, our MFWM can generate the corresponding latent states by turning the embedder’s output, $e_{\textrm{task}}$, into sequences of latent states $s_{t:t+k}$ (examples are shown in Figure 1). The objective of the policy model $\pi_{\theta}$ is then to match the goals specified by the user by performing trajectory matching.
The trajectory matching problem can be solved as a divergence minimization problem [13], between the distribution of the states visited by the policy $\pi_{\theta}$ and the trajectory generated using the aligner-connector networks from the user-specified prompt:
$\theta=\arg\min_{\theta}\mathbb{E}_{a_{t}\sim\pi_{\theta}(s_{t})}\Bigl{[}\sum_{t}\gamma^{t}\texttt{distance}\bigl{(}p_{\phi}(s_{t+1}|s_{t},a_{t})\|p_{\psi}(s_{t+1}|e_{\textrm{task}})\bigr{)}\Bigr{]},\quad\textrm{with}\quad e_{\textrm{task}}=f_{\textrm{PT}}(\cdot).$ (2)
The KL divergence is a natural candidate for the distance function [13]. However, in practice, we found that using the cosine distance between linear projections of the latent states notably speeds up learning and enhances stability. We can then turn the objective in Eq. 2 into a reward for RL:
$r_{\textrm{GenRL{}}}=\cos\bigl{(}g_{\phi}(s^{\textrm{dyn}}_{t+1}),g_{\phi}(s^{\textrm{task}}_{t+1})\bigr{)},\quad\textrm{with}\quad s^{\textrm{dyn}}_{t+1}\sim p_{\phi}(s_{t+1}|s_{t},a_{t}),s^{\textrm{task}}_{t+1}\sim p_{\psi}(s_{t+1}|e_{\textrm{task}}),$ (3)
where $g_{\phi}$ represents the first linear layer of the world model’s decoder. We train an actor-critic model to maximize this reward and achieve the tasks specified by the user [25]. Additional implementation details are provided in Appendix B.
Temporal alignment. One issue with trajectory matching is that it assumes that the distribution of states visited by the agent starts from the same state as the target distribution. However, the initial state generated by the connector may differ from the initial state where the policy is currently in. For example, consider the Stickman agent on the right side of Figure 1. If the agent is lying on the ground and tasked to run, the number of steps to get up and reach running states may surpass the temporal span recognized by the VLM (e.g. typically 4, 8, or 16 frames), causing disalignment in the reward.
To address this initial condition alignment issue, we propose a best matching trajectory technique, inspired by best path decoding in speech recognition [19]. Our technique involves two steps:
  1. 1.
    We compare the first $b$ states of the target trajectory with $b$ states obtained from the trajectories imagined by the agent by sliding along the time axis. This allows one to find at which timestep $t_{a}$ the trajectories are best aligned (the comparison provides the highest reward).
  2. 2.
    We align the temporal sequences in the two possible contexts: (a) if a state from the agent sequence comes before $t_{a}$, the reward uses the target sequence’s initial state; and (b) if the state comes $k$ steps after $t_{a}$, it’s compared to the $s_{t+k}$ state from the target sequence.
In all experiments, we fix $b=8$ (number of frames of the VLM we use [54]), which we found to strike a good compromise between comparing only the initial state ($b=1$) and performing no alignment ($b=\textrm{imagination horizon})$. An ablation study can be found in Appendix E.
4 Experiments
Overall, we employ a set of 4 locomotion environments (Walker, Cheetah, Quadruped, and a newly introduced Stickman environment) [51] and one manipulation environment (Kitchen) [21], for a total of 35 tasks where the agent is trained without rewards, only from text prompts. This represents the first large-scale study of multi-task generalization from language in RL. Details about datasets, tasks, and prompts used can be found in the Appendix C.
4.1 Offline RL
In offline RL, the objective of the agent is to learn to extract a certain task behavior from a given fixed dataset [32]. The performance of the agent generally depends on its ability to ‘retrieve’ the correct behaviors in the dataset and interpolate among them. Popular techniques for offline RL include off-policy RL methods, such as TD3 [17], advantage-weighted behavior cloning, such as IQL [30], and behavior-regularized approaches, such as CQL [31] or TD3+BC [16].
We aim to assess the multi-task capabilities of different approaches for designing rewards using VLMs. We collected large datasets for each of the domains evaluated, containing a mix of structured data (i.e. the replay buffer of agent learning to perform some tasks) and unstructured data (i.e. exploration data collected using [48]). We have removed the explicit reward information about the task and replaced it with a short task description, in language form.
We compare GenRL to two main categories of approaches:
  • Image-language rewards: following [45], the cosine similarity between the embedding for the language prompt and the embedding for the agent’s visual observation is used as a reward. For the VLM, we adopt the SigLIP-B [60] model as it’s reported to have superior performance than the original CLIP [42].
  • Video-language rewards: similar to the image-language rewards, with the difference that the vision embedding is computed from a video of the history of the last $k$ frames, as done in [15]. For the VLM, we use the InternVideo2 model [54], the same used for GenRL.
We test both approaches with a variety of offline RL methods, including IQL, TD3+BC, and TD3. All methods are trained for 500k gradient steps, and evaluated on 20 episodes. Other training details are reported in Appendix D.
Table 1: Behavior extraction. Offline RL from language prompts on tasks that are included in the agent’s training dataset. Scores are episodic rewards averaged over 5 seeds (± standard error) rescaled using min-max scaling with $(\min=\textrm{random policy},\max=\textrm{expert policy})$.
Image-language reward Video-language reward Ours
IQL TD3+BC TD3 IQL TD3+BC TD3 GenRL
walker stand 0.68 ± 0.03 0.97 ± 0.01 0.92 ± 0.06 0.72 ± 0.05 0.59 ± 0.05 1.0 ± 0.0 1.01 ± 0.0
walker run 0.24 ± 0.03 0.27 ± 0.02 0.11 ± 0.03 0.26 ± 0.02 0.25 ± 0.02 0.35 ± 0.01 0.76 ± 0.01
walker walk 0.37 ± 0.04 0.25 ± 0.03 0.15 ± 0.0 0.43 ± 0.04 0.39 ± 0.01 0.86 ± 0.04 0.95 ± 0.03
cheetah run 0.37 ± 0.07 -0.01 ± 0.0 -0.01 ± 0.0 0.15 ± 0.03 -0.01 ± 0.0 0.34 ± 0.01 0.74 ± 0.01
quadruped stand 0.32 ± 0.06 0.4 ± 0.06 0.62 ± 0.04 0.39 ± 0.04 0.41 ± 0.13 0.73 ± 0.06 0.98 ± 0.0
quadruped run 0.3 ± 0.03 0.35 ± 0.01 0.25 ± 0.02 0.36 ± 0.04 0.35 ± 0.04 0.23 ± 0.02 0.61 ± 0.04
quadruped walk 0.16 ± 0.02 0.18 ± 0.02 0.37 ± 0.02 0.19 ± 0.04 0.16 ± 0.04 0.37 ± 0.02 0.73 ± 0.06
stickman run 0.2 ± 0.02 0.25 ± 0.02 0.03 ± 0.0 0.24 ± 0.02 0.18 ± 0.03 0.21 ± 0.0 0.35 ± 0.02
stickman walk 0.43 ± 0.07 0.52 ± 0.05 0.18 ± 0.01 0.47 ± 0.02 0.45 ± 0.08 0.42 ± 0.03 0.78 ± 0.02
stickman stand 0.52 ± 0.05 0.68 ± 0.04 0.56 ± 0.04 0.56 ± 0.06 0.47 ± 0.02 0.61 ± 0.03 0.49 ± 0.04
kitchen microwave 0.14 ± 0.13 0.0 ± 0.0 0.0 ± 0.0 0.02 ± 0.02 0.0 ± 0.0 0.46 ± 0.18 0.67 ± 0.27
kitchen light 0.18 ± 0.12 1.0 ± 0.0 0.66 ± 0.07 0.04 ± 0.02 0.0 ± 0.0 0.0 ± 0.0 0.43 ± 0.19
kitchen burner 0.14 ± 0.07 0.43 ± 0.14 0.02 ± 0.02 0.02 ± 0.02 0.0 ± 0.0 0.28 ± 0.08 0.67 ± 0.14
kitchen slide 0.16 ± 0.1 0.0 ± 0.0 0.0 ± 0.0 0.06 ± 0.04 0.0 ± 0.0 1.0 ± 0.0 1.0 ± 0.0
overall 0.30 ± 0.04 0.38 ± 0.02 0.28 ± 0.02 0.28 ± 0.02 0.23 ± 0.02 0.49 ± 0.04 0.73 ± 0.05
Behavior extraction. We want to verify whether the methods can retrieve the tasks behaviors that are certainly present in the dataset. We present results in Table 1, with episodic rewards rescaled so that 0 represents the performance of a random agent, while 1 represents the performance of an expert agent.
GenRL excels in overall performance across all domains, particularly in dynamic tasks like walking and running in the quadruped and cheetah domains. In contrast, in static tasks such as ’stickman stand’ and kitchen tasks, other methods occasionally outperform GenRL. This can be explained by the fact that the target sequences that GenRL infers from the prompt are often slightly in motion, even in static cases. To address this, we could set the target sequence length to 1 for static prompts, but we opted to maintain the method’s simplicity and generality, acknowledging this as a minor limitation.
As expected, video-language rewards tend to perform better than image-language rewards for dynamic tasks. For video-based rewards, the less conservative approach, TD3, performs better than all other baselines in most tasks, similarly to what is shown in [58]. We instead observe an opposite trend for image-language rewards, where more conservative approaches, such as IQL and TD3+BC, tend to perform better. We believe this is because when the ‘task target’ is static, imitating segments of trajectories from the dataset proves beneficial.
Figure 3: Multi-task generalization. Offline RL from language prompts on tasks that are not deliberately included in the training dataset. Performance averaged over 5 seeds and standard error was reported with black lines. Detailed results per task in Appendix E.
Multi-task generalization. To assess multi-task generalization, we defined a set of tasks not included in the training data. Although we don’t anticipate agents matching the performance of expert models, higher scores in this benchmark help gauge the generalization abilities of different methods. We averaged the performance across various tasks for each domain and summarized the findings in Figure 3, with detailed task results in Appendix E.
Overall, we observe a similar trend as for the behavior extraction results. GenRL significantly outperforms the other approaches, especially in the quadruped and cheetah domains, where the performance is close to the specialized agents’ performance. Both for image-language (-I in the Figure) and video-language (-V in the Figure) more conservative approaches, such as IQL and TD3+BC tend to perform worse. This could be associated with the fact that imitating segments of trajectories is less likely to lead to high-rewarding trajectories, as the tasks are not present in the training data.
4.2 Data-free RL
Figure 4: Data-free RL. Generalization to new tasks, learning behaviors in imagination without relying on any data for initializing the sequences to learn in imagination. Performance is averaged over 5 seeds and standard error is reported with black lines. Detailed results per task in Appendix E.
In the previous section, we evaluated several approaches for designing reward using foundation VLMs. Clearly, model-free RL approaches require continuous access to a dataset, to train the actor-critic and generalize across new tasks. Model-based RL can learn the actor-critic in imagination. However, in previous work [25, 23], imagining sequences for learning behaviors first requires processing actual data sequences. The data is used to initialize the dynamics model, and obtain latent states that represent the starting states to rollout the policy in imagination. Furthermore, in order to learn new tasks, reward-labelled data is necessary to learn a reward model, which provides rewards to the agent during the task learning process.
Foundation models [41] are generally trained on enormous datasets in order to generalize to new tasks. The datasets used for the model pretraining are not necessary for the downstream applications, and sometimes these datasets are not even publicly available [40, 18]. In this section, we aim to establish a new paradigm for foundation models in RL, which follows the same principle of foundation models for vision and language. We call this paradigm data-free RL and we define it as the ability to generalize to new tasks, after pre-training, using no additional data.
GenRL enables data-free RL thanks to two main reasons: the agent learns a task-agnostic MFWM on a large varied dataset during pre-training, and the MFWM enables the possibility of specifying tasks directly in latent space, without requiring any data. Thus, in order to learn behaviors in imagination, the agent can: (i) sample random latent states in the world model’s representation, (ii) rollout sequences in imagination, following the policy, and (iii) compute rewards, using the targets obtained by processing the given prompts with the connector-aligner networks.
Initial states distribution. Uniform sampling from the latent space of the world model often results in meaningless latent states. Additionally, the sequential dynamics model of the MFWM, using a GRU, requires some ’warmup’ steps to discern dynamic environmental attributes, such as velocities.
To address these issues we perform two operations. First, we combine uniformly sampled states from the discrete latent spaces with states generated by randomly sampling the connector model, as sequences generated by the connector tend to have a more coherent structure than random uniform samples. Second, we perform a rollout of five steps using a mix of actions from the trained policy and random actions. This leads to a varied distribution of states, containing dynamic information, which we use as the initial states for the learning in imagination process.
Performance of data-free RL. Our results, detailed in Figure 4, compare data-free RL to traditional offline RL with GenRL, as discussed in Section 4.1. We ablate the choice of using random samples from the connector-aligner to improve randomly sampled initial states. While data-free RL generally shows a slight decrease in overall performance, the differences are minimal across most domains, and it even outperforms in the kitchen domain. The use of states from the connector model enhances average scores and reduces variance, especially noticeable in the cheetah domain.
By employing data-free learning, after pre-training, agents can master new tasks without data, often converging within only 30 minutes of GPU training. As we scale up foundation models for behavior learning, the ability to learn data-free will become crucial. Although very large datasets will be employed to train new foundation models, GenRL adapts well without direct access to original data, offering flexibility where data may be proprietary, licensed or unavailable.
4.3 Training data distribution
Figure 5: Training data distribution. Analysing the impact of the training data distribution on the generalization performance of GenRL. Performance is obtained by training behaviors in data-free mode, after training the MFWM on different subsets of the training dataset. Performance averaged over 3 seeds (black lines indicate standard error). Full results in Appendix E.
As demonstrated in Sections 4.1 and 4.2, after training on a large dataset, a GenRL agent can adapt to multiple new tasks without additional data. The nature of the training data, detailed in Appendix C, combines exploration and task-specific data. To identify critical data types for GenRL, we trained different MFWMs on various dataset subsets. Then, we employ data-free RL to train task behaviors, with analyses over subsets of the walker dataset provided in Figure 5, where ‘all’ reports the data-free performance when training on the full dataset.
The results confirm that a diverse data distribution is crucial for task success, with the best performance achieved by using the complete dataset, followed by the varied exploration data. Task-specific data effectiveness depends on task complexity, for instance, ’run’ data proves more generalizable than ’walk’ or ’stand’ data across tasks. Crucially, ’stand’ data, which shows minimal variation, limits learning for a general agent but can still manage simpler tasks like ’lying down’ and ’sitting on knees’ as detailed in Appendix E.
Moving forward with training foundation models in RL, it will be essential to develop methods that extract multiple behaviors from unstructured data and accurately handle complex behaviors from large datasets. Thus, the ability of GenRL to primarily leverage unstructured data is a significant advantage for scalability.
5 Additional Analysis
A framework for behavior generation. A common challenge with using LLMs and VLMs involves the need for prompt tuning to achieve specific tasks. As GenRL relies on a foundation VLM, similar to previous approaches [4, 45] it is not immune from this issue. However, GenRL uniquely allows for the visualization of targets obtained from specific prompts. By decoding the latent targets, using the MFWM decoder, we can visualize the interpreted prompt before training the corresponding behavior. This enables a much more explainable framework, which allows fast iteration for prompt tuning, compared to previous (model-free) approaches which often require training the agent to identify which behaviors are rewarded given a certain prompt.
Video grounding. In Section 4, we have focussed our evaluation on language-driven RL, using text prompts to specify tasks. While language strongly simplifies specification of the task, in some cases providing visual examples of the task might be easier. Similarly as for language prompts, GenRL allows translating vision prompts (short videos) into behaviors. As this is a less common evaluation setting, we limit our assessment to qualitative results.
In Figure 6(a), we present a set of video grounding examples, obtained by inferring the latent targets corresponding to the vision prompts (right image) and then using the decoder model to decode images (left image). We observe that the agent is able to translate short human action videos [^2] Short videos are generated using meta.ai. into the same actions but for the Stickman embodiment. By applying this approach, it would be possible to learn behaviors from a single video.
Scaling to complex observations. Generalist embodied agents should be able to scale to open-ended learning settings. Using GenRL, we explored this by training an agent in the Minecraft environment using a small dataset collected by a DreamerV3 agent [25]. The primary challenge we found was the model’s difficulty in reconstructing complex observations in this open-ended environment.
Reconstructing complex observations is a common issue with world models [10]. To overcome this limitation, while keeping the method unaltered, we attempted to scale up the number of parameters of MFWM. Qualitative reconstruction results are presented in Figure 6(b). We observe that the agent is able to identify different biomes from language, even with the smaller size of the model. However, the reconstructions are significantly blurrier compared to the other environments we analyzed (e.g. Figure 6(a)). When using a larger model, the reconstructions gain some details but the results still highlight the difficulty of the model in providing accurate targets from prompts.
While this might not be an issue for simple high-level tasks, e.g. ‘navigate to a beach’, unclear targets might make it difficult to perform more precise actions, e.g. ‘attack a zombie’. Future research should aim to address this issue, for instance, by improving our simple GRU-based architecture, leveraging transformers or diffusion models to improve the quality of the representation [29, 3].
(a) Video grounding examples
(b) Decoded language prompts in Minecraft
6 Discussion
We introduced GenRL, a world-model based approach for grounding vision-language prompts into embodied domains and learning the corresponding behaviors in imagination. The multimodal foundation world models of GenRL can be trained using unimodal data, overcoming the lack of multimodal data in embodied RL domains. The data-free RL property of GenRL lays the groundwork for foundation models in RL that can generalize to new tasks without additional data.
Limitations. Despite its strengths, GenRL presents some limitations, largely due to inherent weaknesses in its components. From the VLMs, GenRL inherits the issue related to the multimodality gap [33, 61] and the reliance on prompt tuning. We proposed a connection-alignment mechanism to mitigate the former. For the latter, we presented an explainable framework, which facilitates prompt tuning by allowing decoding of the latent targets corresponding to the prompts. From the world model, GenRL inherits a dependency on reconstructions, which offers advantages such as explainability but also drawbacks, such as failure modes with complex observations.
Future work. As we strive to develop foundation models for generalist embodied agents, our framework opens up numerous research opportunities. One such possibility is to learn multiple behaviors and have another module, e.g. an LLM, compose them to solve long-horizon tasks. Another promising area of research is investigating the temporal flexibility of the GenRL framework. We witnessed that for static tasks, greater temporal awareness could enhance performance. This concept could also apply to actions that extend beyond the time comprehension of the VLM. Developing general solutions to these challenges could lead to significant advancements in the framework.
Acknowledgments
Pietro Mazzaglia is funded by a Ph.D. grant of the Flanders Research Foundation (FWO). This research was supported by a Mitacs Accelerate Grant.
Appendix A Extended Related Work
World models. Recent research has focused on the question of how to learn world models from large-scale video datasets [36, 56]. In [6], they leverage a latent action representation, but their work is mostly focussed on 2D platform videogames or simple robotic actions. In [14], they use frame-by-frame video prediction as a way to provide rewards for RL. DynaLang [35] studies the incorporation of language prediction as part of the world model, to train multimodal world models also from datasets without actions or rewards. The representation in DynaLang is shared in the world model between vision and language, while for GenRL, the world model representation is trained on vision-only data and connected-aligned to the multimodal foundation representation.
Foundation models for actions. Few cases of foundation models for embodied domains have been developed. Notable mentions are GATO [44], a large-scale behavior cloning agent, trained on 604 tasks. VPT [2] a large-scale model trained on Minecraft data, using human-expert labeled trajectories. The model learns strong behavioral priors by behavior cloning which can be fine-tuned using RL. STEVE-1 [34] connects VPT’s behavioral prior with the MineCLIP model representation [15], using the unCLIP approach [43]. RT-X [12] are large-scale trasnformer models trained on expert robotics dataset, sharing a common action space (end-effector pose) across different embodiments.
Appendix B Implementation details
Actor-critic. Rewards can be maximized over time in imagination in a RL fashion, using actor-critic models of the form:
$\begin{split}\textrm{Actor:}\quad\pi_{\theta}(a_{t}|s_{t}),\end{split}\quad\quad\begin{split}\textrm{Critic:}\quad v_{\theta}(R^{\lambda}_{t}|s_{t}),\quad\textrm{where}\quad R^{\lambda}_{t}=r_{t}+\gamma[(1-\lambda v_{t+1})+\lambda R^{\lambda}_{t+1}]\end{split}$
For the actor-critic, we follow the implementation advances proposed in DreamerV3 [25] (version 1 of the paper, dated January 2023), such as using a two-hot distribution for learning the critic network and scaling returns in the actor loss.
When computing the reward $r_{\textrm{GenRL}}$, we use the mode of the distribution for the target $s^{\textrm{task}}_{t+1}\sim p_{\psi}(s_{t+1}|e_{\textrm{task}})$ to improve stability.
Hyperparameters. For the hyperparameters, we follow DreamerV3 [25] (version 1 of the paper, dated January 2023). Differences from the default hyperparameters or model size choices are illustrated in Table 2. For instance, a main difference is that we use difference batch sizes/lengths for training the MFWM and the actor-critic as these two stages are now independent from each other.
The connector network uses the same hyperparameters and architecture as the sequential dynamics of the world model. The aligner network employs a small U-Net, with a bottleneck that is half the size of the embedding representation.
Name Value
Multimodal Foundation World Model
Batch size 48
Sequence length 48
GRU recurrent units 1024
CNN multiplier 48
Dense hidden units 1024
MLP layers 4
Actor-Critic
Batch size 32
Sequence length 32
Table 2: World model and actor-critic hyperparameters.
Appendix C Tasks
We present the list of tasks employed, along with the language prompts used for specifying the task, in Table 3. We introduce a new embodied environment, the ‘Stickman’, which serves as a humanoid robot (compared to the one present in the dm_control suite) that is simpler to control, thanks to a reduced number of joints. Its addition allows studying behaviors that involve upper body movements, rather than focusing on lower body motions. For the newly introduced tasks, the goal can be easily inferred by reading the task’s name or its prompt. For the ‘flipping’ tasks, we consider flips both in forward direction and backward direction, as the VLM struggles to distinguish directions.
Table 3: Task and prompt used for each task
Task Prompt Specialized agent Random agent
score score
quadruped run spider running fast 930 10
quadruped walk spider walking fast 960 10
quadruped stand spider standing 990 15
quadruped jump spider jumping 875 15
quadruped two legs on two legs 875 14
quadruped lie down lying down 965 750
cheetah run running like a quadruped 890 9
cheetah standing standing like a human 930 5
cheetah lying down lying down 920 430
stickman walk robot walk fast clean 960 35
stickman run robot run fast clean 830 25
stickman stand standing 970 70
stickman flipping doing flips 790 45
stickman one foot stand on one foot 865 20
stickman high kick stand up and kick 920 55
stickman lying down lying down horizontally 965 380
stickman sit knees praying 966 40
stickman lunge pose lunge pose 950 100
stickman headstand headstand 955 180
stickman boxing punch 920 80
stickman hands up standing with the hands up 830 5
walker walk walk fast clean 960 45
walker run run fast clean 770 30
walker stand standing up straight 970 150
walker flipping doing backflips 720 20
walker one foot stand on one foot 955 20
walker high kick stand up and kick 960 25
walker lying down lying down horizontally 975 170
walker sit knees praying 945 100
walker lunge pose lunge pose 945 150
kitchen microwave opening the microwave fully open 1 0
kitchen light activate the light 1 0
kitchen burner the burner becomes red 1 0
kitchen slide slide cabinet above the knobs 1 0
The prompts we use have been fine-tuned for the InternVideo2 model [54]. However, we found that they mostly improved performance for the SigLIP model too [60]. One common observation is that these models are generally biased towards human actions. Thus, specifying the embodiment in the prompt is sometimes helpful, e.g. ‘spider running fast’ or ‘running like a quadruped’. Another observation is that for some behaviors the agent can produce very different styles, e.g. the agent can be walking in a slow or fast way, or in a more or less composed manner. Specifying words like ‘fast’ or ‘clean’ helps clarifying what kind of behavior is expected.
Appendix D Experiments settings
Baselines. In order to implement performant offline RL baselines we adopt the findings of [49] and [7], adopting larger deeper networks and layer normalization.
Inputs are 64x64x3 RGB images. We use a frame stack of 3. The encoder architecture is adapted from the DrQ-v2 encoder [59]. We did find augmentations on the images, e.g. random shifts, to hurt performance.
Offline RL. For each task, training model-free agents (IQL, TD3, TD3+BC) requires re-training the full agent (visual encoder, actor, critic) on the entire dataset, from scratch, while training model-based agents (GenRL) requires training the model once for each domain and then training an actor-critic for each task. Moreover, for training the actor-critic in GenRL, we only use 50k gradient steps, as the policy converges significantly faster than for the other methods.
Compute resources. We use a cluster of V100 with 16GB of VRAM for all our experiments. To enable efficient training, image and video-CLIP embeddings are computed in advance and stored with the datasets. Training the MFWM for 500k gradient steps takes $\sim$ 5 days. After pre-training the MFWM, training the actor-critic for a prompt for 50k gradient steps takes less than 5 hours. In data-free mode, it takes less than 3 hours. In both cases, convergence normally arrives after 10k gradient steps, but we keep training. Model-free baselines take around 7 hours to train for 500k gradient steps. normally arrive
Datasets composition. We present the datasets’ composition in Table 4.
Table 4: Datasets composition.
Domain $\sim$ num of observations Subset Subcount
walker 2.5M walker run 500k
walker walk 500k
walker stand 500k
walker expl 1M
cheetah 1.8M cheetah run 1M
cheetah expl 820k
quadruped 2.5M quadruped expl 1M
quadruped run 500k
quadruped stand 500k
quadruped walk 500k
kitchen 3.6M kitchen slide 700k
kitchen light 700k
kitchen bottom burner 700k
kitchen microwave 700k
kitchen expl 800k
stickman 2.5M stickman stand 500k
stickman walk 500k
stickman expl 1M
stickman run 500k
minecraft 4M - -
Appendix E Additional results
Figure 7: Temporal alignment ablation. We analyze the impact of temporal alignment in our proposed RL objective for matching sequential targets. Results averaged over 3 seeds.
Figure 8: Multi-task generalization detailed results. Results averaged over 5 seeds.
Figure 9: Data-free RL detailed results. Results averaged over 5 seeds.
Figure 10: Training data distribution detailed results. Results averaged over 3 seeds.